# Charged rotating black holes in scalar multipolar universes

### v1

@@ Abstract
Recently, Cardoso and Natário [6] constructed an exact solution of Einstein-scalar field equations that describes a scalar counterpart of the Schwarzschild-Melvin Universe. In fact, this solution belongs to a more general class of solutions described by Herdeiro in [7]. In this work we show how to further generalize these solutions in presence of rotation and various charges. More specifically, we describe general charged rotating black holes with NUT charge and present some of their properties.
@@ 1 Introduction
The purpose of this work is to further generalize the Herdeiro solutions in presence of rotation and also electromagnetic and NUT charges. More precisely, in the next section one shows how to simply obtain such a solution in presence of a scalar field. Our approach is based on the symmetries of the dimensionally reduced theory down to three dimensions. Rotating black holes of this kind could also be constructed using the methods described in [8], however, in our approach one could easily include the effect of electric and magnetic charges as well.
The structure of this paper is as follows: in the next section one describes the solution generating technique. In section $3$ one presents the scalar generalization of the Kerr-Newman black hole with NUT charge in scalar multipolar Universes. We focus on a special solution describing the rotating generalization of the Cardoso-Naváro solution and present some of its properties. The final section is dedicated to conclusions and avenues for further work.
@@ 2 The solution generating technique
$ds^{2}$ $=$ $-e^{-2\psi}dt^{2}+e^{2\psi}\bigg{[}e^{2\gamma}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (1)
@@
$ds^{2}$ $=$ $-e^{-2\psi}dt^{2}+e^{2\psi}\bigg{[}e^{2\lambda}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (14)
@@
$ds^{2}$ $=$ $-e^{-2\psi}dt^{2}+e^{2\psi}\bigg{[}e^{2\gamma+2\mu}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]}.$ (18)
This is the class of solutions considered by Herdeiro in [7]. In the followings we shall show that it is easy to generalize this construction in the presence of rotation and various other charges.
@@
$ds^{2}$ $=$ $-e^{2\psi}(dt+{\cal A})^{2}+e^{2\psi}\bigg{[}e^{2\gamma}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (19)
@@
${\cal L}_{3}$ $=$ $\sqrt{g_{3}}\bigg{[}R_{3}-2(\partial\psi)^{2}+e^{-4\psi}({\cal F})^{2}-e^{-2\psi}F_{(2)}^{2}+2e^{2\psi}(\partial\chi)^{2}\bigg{]},$ (21)
while the three-geometry is given by (2). Here we defined $F_{(2)}=dA_{\varphi}\wedge d\varphi-d\chi\wedge{\cal A}$ and ${\cal F}=d{\cal A}$. One can write the three-dimensional Einstein equations in the following form:
$R_{ij}[\gamma]$ $=$ $T_{ij}^{(1)}=2(\partial_{i}\psi)(\partial_{j}\psi)-2e^{-4\psi}\bigg{[}{\cal F}_{ik}{\cal F}_{j}^{~{}k}-\frac{1}{2}g_{ij}{\cal F}_{kl}{\cal F}^{kl}\bigg{]}-2e^{2\psi}(\partial_{i}\chi)(\partial_{j}\chi)+$ (22)
$+$ $2e^{-2\psi}\bigg{[}F_{(2)ik}F_{(2)j}^{~{}k}-\frac{1}{2}g_{ij}F_{(2)kl}F_{(2)}^{kl}\bigg{]}.$
@@
$\nabla_{i}\left(e^{-4\psi}{\cal F}^{ij}\right)$ $=$ $0,~{}~{}~{}~{}~{}\nabla_{i}\left(e^{-2\psi}F_{(2)}^{ij}\right)=0,~{}~{}~{}~{}~{}\nabla_{i}\left(e^{2\psi}\nabla^{i}\chi\right)=0,$
$\Delta\psi$ $=$ $-\frac{1}{2}e^{-2\psi}F_{(2)}^{2}+e^{-4\psi}{\cal F}^{2}+e^{2\psi}(\partial\chi)^{2}.$ (23)
@@
In general, one can write it in the Weyl-Papapetrou form:
$ds^{2}$ $=$ $-e^{2\psi}(dt+{\cal A})^{2}+e^{2\psi}\bigg{[}e^{2\lambda}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]}.$ (25)
@@
${\cal L}_{3}$ $=$ $\sqrt{g_{3}}\bigg{[}R_{3}-2(\partial\psi)^{2}+e^{-4\psi}({\cal F})^{2}-e^{-2\psi}F_{(2)}^{2}+2e^{2\psi}(\partial\chi)^{2}-2(\partial_{\mu}\phi)(\partial^{\mu}\phi)\bigg{]}.$ (26)
@@
$R_{ij}[\lambda]$ $=$ $T^{(1)}_{ij}+T^{2}_{ij},$ (27)
where $T^{(1)}_{ij}=R_{ij}[\gamma]$ is listed in (22) while $T^{(2)}_{ij}=2(\partial_{i}\phi)(\partial_{j}\phi)=R_{ij}[\mu]$. Thence one obtains $R_{ij}[\lambda]=R_{ij}[\gamma]+R_{ij}[\mu]$ and therefore one can simply take $\lambda=\gamma+\mu$ to simply satisfy Einstein’s field equations in three dimensions.
@@
In conclusion, starting with a known solution (19) of the Einstein-Maxwell equations then the new solution of the field equations derived from the Lagrangean (24) can be written in the form:
$ds^{2}$ $=$ $-e^{2\psi}(dt+{\cal A})^{2}+e^{2\psi}\bigg{[}e^{2\gamma+2\mu}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (28)
where the scalar field takes the form (12), while the function $\mu$ is given by (13).
3 The Kerr-Newman-NUT solution with a scalar field
As an example of the solution generating technique described in the previous section, let us consider the Kerr-Newman-NUT solution in four dimensions (see for instance [10], [11]) [^2] Note that we could also consider the most general solution with the acceleration parameter, however, for simplicity we set it to zero here.:
$ds^{2}$ $=$ $-\frac{\Delta_{r}}{\Sigma}\bigg{[}dt-\left(a\sin^{2}\theta+2l(a-\cos\theta)\right)d\varphi\bigg{]}^{2}+\Sigma\left(\frac{dr^{2}}{\Delta_{r}}+d\theta\right)^{2}+$ (29)
$+$ $\frac{\sin^{2}\theta}{\Sigma}\big{[}adt-\left(r^{2}+(a+l)^{2}\right)d\varphi\big{]}^{2},$
@@ 3 The Kerr-Newman-NUT solution with a scalar field
$\Delta_{r}$ $=$ $r^{2}-2mr+a^{2}+e^{2}+g^{2}-l^{2},~{}~{}~{}\Sigma=r^{2}+(l+a\cos\theta)^{2}.$ (30)
Here $m$ is the mass parameter, $a$ is the rotation parameter, $l$ is the NUT charge, while $e$ and $g$ are the electric and magnetic charges. The components of the electromagnetic Maxwell field $A=\chi dt+A_{\varphi}d\varphi$ are:
$\chi$ $=$ $-\frac{er+g(l+a\cos\theta)}{\Sigma},$
$A_{\varphi}$ $=$ $\frac{er\big{[}(a+l)^{2}-(l+a\cos\theta)^{2}\big{]}+g(1+a\cos\theta)\big{[}r^{2}+(a+l)^{2}\big{]}}{a\Sigma}.$ (31)
As it is well-known, this geometry can be brought into the Weyl-Papapetrou form (19) by using the following coordinate transformations:
@@
$d\rho^{2}+dz^{2}$ $=$ $(\Delta_{r}\cos^{2}\theta+(r-m)\sin^{2}\theta)\left(\frac{dr^{2}}{\Delta_{r}}+d\theta^{2}\right),$ (33)
@@
$e^{2\gamma}$ $=$ $\frac{\Delta_{r}-a^{2}\sin^{2}\theta}{\Delta_{r}\cos^{2}\theta+(r-m)\sin^{2}\theta}.$ (34)
@@
$ds^{2}$ $=$ $-\frac{\Delta_{r}}{\Sigma}\bigg{[}dt-\left(a\sin^{2}\theta+2l(a-\cos\theta)\right)d\varphi\bigg{]}^{2}+\Sigma e^{2\mu}\left(\frac{dr^{2}}{\Delta_{r}}+d\theta\right)^{2}+$ (35)
$+$ $\frac{\sin^{2}\theta}{\Sigma}\big{[}adt-\left(r^{2}+(a+l)^{2}\right)d\varphi\big{]}^{2},$
where the Maxwell field components are given in (31), the scalar field $\phi$ has the general expression (12), while the function $\mu$ is given by (13), using (32).
@@
$\phi$ $=$ $kz=k(r-m)\cos\theta,~{}~{}~{}~{}~{}e^{2\mu}=e^{-k^{2}\rho^{2}}=e^{-k^{2}\Delta_{r}\sin^{2}\theta}.$ (38)
If one sets to zero all the charges one obtains the rotating version of the scalar-Schwarzschild-Melvin [6] in the form:
@@
where $\Delta_{r}=r^{2}-2mr+a^{2}$ and $\Sigma=r^{2}+a^{2}\cos^{2}\theta$. It is easy to check that it satisfies the Einstein-scalar equations $R_{\mu\nu}=2(\partial_{\mu}\phi)(\partial_{\nu}\phi)$. If $a=0$ it reduces to the solution in [6]. Its Ricci scalar can be simply expressed as:

### v4

# Accelerating charged and rotating black holes in scalar multipolar universes

@@ Abstract
Recently, Cardoso and Natário [6] constructed an exact solution of Einstein-scalar field equations that describes a scalar counterpart of the Schwarzschild-Melvin Universe. In fact, this solution belongs to a more general class of solutions described by Herdeiro in [7]. In this work we show how to further generalize these solutions in presence of acceleration, rotation and various charges. More specifically, we describe general accelerating charged rotating black holes with NUT charge in scalar multipolar universes and present some of their properties.
@@ 1 Introduction
The purpose of this work is to further generalize the Herdeiro solutions in presence of acceleration, rotation and also electromagnetic and NUT charges. More precisely, in the next section one shows how to simply obtain such a solution in presence of a scalar field. Our approach is based on the symmetries of the dimensionally reduced theory down to three dimensions. Rotating black holes of this kind could also be constructed using the methods described in [8], however, in our approach one could easily include the effect of electric and magnetic charges as well. One should note that the subject of exact solutions of Einstein-scalar field equations in four and higher dimensions has a long history: in fact, in section $2$ we will recover a previously known result discovered first in [9]. Based on this we were able to easily construct the most general charged and rotating black holes with minimally coupled and massless scalar fields, a solution encompassing various other accelerating, charged and/or rotating solutions known in literature [10], [11], [12], [13]. One should however note that our approach is based on the use of Weyl-Papapetrou coordinates and as such it can only describe black holes with scalar fields up to five dimensions. There are however known methods to add scalar fields to static geometries in higher than five dimensions (see for instance [14] and references therein). Presumably, this kind of solutions could be generalized to higher than five dimensions using the methods of [11].
The structure of this paper is as follows: in the next section one describes the solution generating technique. In section $3$ one presents the scalar generalization of the Type D accelerating Kerr-Newman black hole with NUT charge in scalar multipolar Universes. We focus on a special solution describing the accelerating and rotating generalization of the Cardoso-Naváro solution and present some of its properties. The final section is dedicated to conclusions and avenues for further work.
@@ 2 The solution generating technique
$ds^{2}$ $=$ $-e^{2\psi}dt^{2}+e^{-2\psi}\bigg{[}e^{2\gamma}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (1)
@@
$ds^{2}$ $=$ $-e^{2\psi}dt^{2}+e^{-2\psi}\bigg{[}e^{2\lambda}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (14)
@@
$ds^{2}$ $=$ $-e^{2\psi}dt^{2}+e^{-2\psi}\bigg{[}e^{2\gamma+2\mu}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (18)
where the scalar field $\phi$ is given by (12) and the function $\mu$ is given by (13). This is the class of solutions considered by Herdeiro in [7]. In the followings we shall show that it is easy to generalize this construction in the presence of rotation and various other charges.
@@
$ds^{2}$ $=$ $-e^{2\psi}(dt+{\cal A})^{2}+e^{-2\psi}\bigg{[}e^{2\gamma}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (19)
@@
${\cal L}_{3}$ $=$ $\sqrt{g_{3}}\bigg{[}R_{3}-2(\partial\psi)^{2}+e^{4\psi}({\cal F})^{2}-e^{2\psi}F_{(2)}^{2}+2e^{-2\psi}(\partial\chi)^{2}\bigg{]},$ (21)
while the three-geometry is given by (2). Here we defined $F_{(2)}=dA_{\varphi}\wedge d\varphi-d\chi\wedge{\cal A}$ and ${\cal F}=d{\cal A}$, while we denoted $(\partial\psi)^{2}=(\partial_{i}\psi)(\partial^{i}\psi)$, ${\cal F}^{2}={\cal F}_{ij}{\cal F}^{ij}$, $F_{(2)}^{2}=F_{(2)ij}F_{(2)}^{ij}$ and $(\partial\chi)^{2}=(\partial_{i}\chi)(\partial^{i}\chi)$. One can write the three-dimensional Einstein equations in the following form:
$R_{ij}[\gamma]$ $=$ $T_{ij}^{(1)}=2(\partial_{i}\psi)(\partial_{j}\psi)-2e^{4\psi}\bigg{[}{\cal F}_{ik}{\cal F}_{j}^{~k}-\frac{1}{2}g_{ij}{\cal F}_{kl}{\cal F}^{kl}\bigg{]}-2e^{-2\psi}(\partial_{i}\chi)(\partial_{j}\chi)+$ (22)
$+$ $2e^{2\psi}\bigg{[}F_{(2)ik}F_{(2)j}^{~k}-\frac{1}{2}g_{ij}F_{(2)kl}F_{(2)}^{kl}\bigg{]}.$
@@
$\nabla_{i}\left(e^{4\psi}{\cal F}^{ij}\right)$ $=$ $0,~~~~~\nabla_{i}\left(e^{2\psi}F_{(2)}^{ij}\right)=0,~~~~~\nabla_{i}\left(e^{-2\psi}\nabla^{i}\chi\right)=0,$
$\Delta\psi$ $=$ $\frac{1}{2}e^{2\psi}F_{(2)}^{2}-e^{4\psi}{\cal F}^{2}-e^{-2\psi}(\partial\chi)^{2}.$ (23)
@@
In general, let us assume that one can write it in the Weyl-Papapetrou form:
$ds^{2}$ $=$ $-e^{2\psi}(dt+{\cal A})^{2}+e^{-2\psi}\bigg{[}e^{2\lambda}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]}.$ (25)
@@
${\cal L}_{3}$ $=$ $\sqrt{g_{3}}\bigg{[}R_{3}-2(\partial\psi)^{2}+e^{4\psi}({\cal F})^{2}-e^{2\psi}F_{(2)}^{2}+2e^{-2\psi}(\partial\chi)^{2}-2(\partial_{i}\phi)(\partial^{i}\phi)\bigg{]}.$ (26)
@@
$R_{ij}[\lambda]$ $=$ $T^{(1)}_{ij}+T^{(2)}_{ij}$, (27)
where $T^{(1)}_{ij}=R_{ij}[\gamma]$ is listed in (22), while $T^{(2)}_{ij}=2(\partial_{i}\phi)(\partial_{j}\phi)=R_{ij}[\mu]$. Thence one obtains $R_{ij}[\lambda]=R_{ij}[\gamma]+R_{ij}[\mu]$ and therefore one can simply take $\lambda=\gamma+\mu$ to simply satisfy Einstein’s field equations in three dimensions.
@@
In conclusion, starting with a known stationary solution (19) of the Einstein-Maxwell equations then the new solution of the field equations derived from the Lagrangean (24) can be written in the form:
$ds^{2}$ $=$ $-e^{2\psi}(dt+{\cal A})^{2}+e^{-2\psi}\bigg{[}e^{2\gamma+2\mu}(d\rho^{2}+dz^{2})+\rho^{2}d\varphi^{2}\bigg{]},$ (28)
where the scalar field $\phi$ takes the general form (12), while the function $\mu$ is given by (13). A similar theorem was previously proven by other means in [9].
3 The accelerating Kerr-Newman-NUT solution with a multipolar scalar field
As an example of the solution generating technique described in the previous section, let us consider the newly discovered Type D accelerating Kerr-Newman-NUT solution in four dimensions (see for instance [16], [17]) [^2] Here we have chosen the value of the constant $C_{f}=\frac{1}{(1+\alpha^{2}a^{2})}$.:
$ds^{2}$ $=$ $\frac{1}{\Omega^{2}}\bigg{[}-\frac{\Delta_{r}}{\Sigma}\left(Adt-Bd\varphi\right)^{2}+\frac{\Sigma}{(1+\alpha^{2}a^{2})}\left(\frac{dr^{2}}{\Delta_{r}}+\frac{dx}{\Delta_{x}}\right)^{2}+$ (29)
$+$ $\frac{\Delta_{x}}{\Sigma}\left(Cdt+Dd\varphi\right)^{2}\bigg{]},$
@@ 3 The accelerating Kerr-Newman-NUT solution with a multipolar scalar field
$\Omega$ $=$ $1-\alpha rx,~~~A=1+\alpha^{2}(l^{2}-a^{2})x^{2},~~~B=a+2lx+ax^{2},$
$C$ $=$ $a+2\alpha lr+\alpha^{2}ar^{2},~~~D=(l^{2}-a^{2})+r^{2}$
$\Delta_{r}$ $=$ $(1-\alpha^{2}r^{2})\big{[}(r-m)^{2}-\sigma^{2}\big{]},~~~\Delta_{x}=(1-x^{2})\big{[}(1-\alpha mx)^{2}-\alpha^{2}x^{2}\sigma^{2}\big{]},$
$\Sigma$ $=$ $(l+ax)^{2}+2\alpha l(a+2lx+ax^{2})r+\alpha^{2}(a^{2}-l^{2})^{2}x^{2}+\big{[}1+\alpha^{2}(a+lx)^{2}\big{]}r^{2},$
$\sigma$ $=$ $\sqrt{m^{2}+l^{2}-a^{2}-e^{2}-g^{2}}.$ (30)
Here $\alpha$ is the acceleration parameter, $m$ is the mass parameter, $a$ is the rotation parameter, $l$ is the NUT charge, while $e$ and $g$ are the electric and magnetic charges. The components of the electromagnetic Maxwell field $A=\chi dt+A_{\varphi}d\varphi$ are listed in [16], [17], [18] and, for simplicity, we will not list them here since the addition of a scalar field in this background has no effect on the Maxwell field components. In this geometry one makes use of the radial coordinate $r$ and the coordinate $x$ that can be related to the usual polar angle $\theta$ by using the relation $x=\cos\theta$. If the acceleration parameter is set equal to zero, $\alpha=0$, one recovers the Kerr-Newman-NUT black hole solution.
As it is well-known, the Kerr-Newman-NUT metric element can be brought into the Weyl-Papapetrou form (19) by using the following coordinate transformations:
@@
$d\rho^{2}+dz^{2}$ $=$ $(\Delta_{r}\cos^{2}\theta+(r-m)^{2}\sin^{2}\theta)\left(\frac{dr^{2}}{\Delta_{r}}+d\theta^{2}\right),$ (32)
@@
$e^{2\gamma}$ $=$ $\frac{\Delta_{r}-a^{2}\sin^{2}\theta}{\Delta_{r}\cos^{2}\theta+(r-m)^{2}\sin^{2}\theta}.$ (33)
@@
$ds^{2}$ $=$ $-\frac{\Delta_{r}}{\Sigma}\bigg{[}dt-\left(a\cos^{2}\theta+2l\cos\theta+a)\right)d\varphi\bigg{]}^{2}+\Sigma e^{2\mu}\left(\frac{dr^{2}}{\Delta_{r}}+d\theta\right)^{2}+$ (34)
$+$ $\frac{\sin^{2}\theta}{\Sigma}\big{[}adt-\left(r^{2}+(l^{2}-a^{2})\right)d\varphi\big{]}^{2},$
where the Maxwell field components are listed in [16], [17], [18], the scalar field $\phi$ has the general expression (12), while the function $\mu$ is given by (13), using (31). Harmonic scalars with growing multipoles (described by terms proportional to $b_{l}$ in the multipolar expansion (12)) lead to naked singularities at spatial infinity, while harmonic scalars with decaying multipoles (described by the terms proportional to $a_{l}$ in (12)) lead to asymptotic flat geometries with singular horizons, in accordance to the no-hair theory.
@@
Let us now consider the presence of the acceleration parameter $\alpha$. In this case it can be shown that the general Type D accelerating Kerr-Newman-NUT solution (29) can still be brought into the Weyl-Papapetrou form by employing the following coordinate transformations:
$\rho$ $=$ $\frac{\sqrt{\Delta_{r}\Delta_{x}}}{(1+\alpha rx)^{2}},~~~z=\frac{(\alpha r+x)\big{[}(r-m)(1+\alpha mx)+\alpha x\sigma^{2}\big{]}}{(1+\alpha rx)^{2}}.$ (37)
Therefore, the general Type D accelerating Kerr-Newman-NUT solution in a scalar multipolar universe can be written in the following symmetric form:
$ds^{2}$ $=$ $\frac{1}{\Omega^{2}}\bigg{[}-\frac{\Delta_{r}}{\Sigma}\left(Adt-Bd\varphi\right)^{2}+\frac{\Sigma}{(1+\alpha^{2}a^{2})}e^{2\mu}\left(\frac{dr^{2}}{\Delta_{r}}+\frac{dx}{\Delta_{x}}\right)^{2}+$ (38)
$+$ $\frac{\Delta_{x}}{\Sigma}\left(Cdt+Dd\varphi\right)^{2}\bigg{]},$
where once again the scalar field $\phi$ is given in (12), while the function $\mu$ is given by (13), using now the coordinate transformations given in (37). The horizons are again determined by the roots of the equation $\Delta_{r}=0$. One obtains the inner horizon, located at $r_{-}=m-\sqrt{\sigma}$, the black hole horizon at $r_{+}=m+\sqrt{\sigma}$, while acceleration horizon is located at $r_{\alpha}=\frac{1}{\alpha}$.
@@
$\phi$ $=$ $kz=k\frac{(\alpha r+x)\big{[}(r-m)(1+\alpha mx)+\alpha x\sigma^{2}\big{]}}{(1+\alpha rx)^{2}},~~~~~e^{2\mu}=e^{-k^{2}\rho^{2}}=e^{-\frac{k^{2}\Delta_{r}\Delta_{x}}{(1+\alpha rx)^{4}}}.$ (39)
If one sets to zero all the charges and the acceleration parameter $\alpha=0$ one obtains $\sigma^{2}=m^{2}-a^{2}$, which corresponds to the rotating version of the scalar-Schwarzschild-Melvin [6] in the form previously found in [19]:
@@
where $\Delta_{r}=r^{2}-2mr+a^{2}$ and $\Sigma=r^{2}+a^{2}\cos^{2}\theta$. It is easy to check that it satisfies the Einstein-scalar equations $R_{\mu\nu}=2(\partial_{\mu}\phi)(\partial_{\nu}\phi)$. If $a=0$ it further reduces to the solution in [6]. Its Ricci scalar can be simply expressed as: